Testing Frankenstein’s parts

version 0.3

                Matt Swift <swift@bu.edu>

1995/02/26

1 New ways to include

This will take several pages.

First, check that we can \include as usual, so the rest of this page should be blank, and the Melville quote should be alone on the next page.

This text is in a separate file that just contains text, and some L^{A}TEX code. For a change, let’s quote Moby Dick, chapter 94:

Squeeze! squeeze! squeeze! all the morning long; I squeezed that sperm till I myself almost melted into it; I squeezed that sperm till a strange sort of insanity came over me; and I found myself unwittingly squeezing my co-laborers’ hands in it, mistaking their hands for the gentle globules. Such an abounding, affectionate, friendly, loving feeling did this avocation beget; that at last I was continually squeezing their hands, and looking up into their eyes sentimentally; as much as to say,—Oh! my dear fellow beings, why should we longer cherish any social acerbities, or know the slightest ill-humor or envy! Come; let us squeeze hands all round; nay, let us all squeeze ourselves into each other; let us squeeze ourselves universally into the very milk and sperm of kindness.

Would that I could keep squeezing that sperm for ever! For now, since by many prolonged, repeated experiences, I have perceived that in all cases man must eventually lower, or at least shift, his conceit of attainable felicity; not placing it anywhere in the intellect or the fancy; but in the wife, the heart, the bed, the table, the saddle, the fire-side, the country; now that I have perceived all this, I am ready to squeeze case eternally. In thoughts of the visions of the night, I saw long rows of angels in paradise, each with his hands in a jar of spermaceti.

This should be on a new page. Now we include the same thing, but it should come immediately here because we use \include*:

This text is in a separate file that just contains text, and some L^{A}TEX code. For a change, let’s quote Moby Dick, chapter 94:

Squeeze! squeeze! squeeze! all the morning long; I squeezed that sperm till I myself almost melted into it; I squeezed that sperm till a strange sort of insanity came over me; and I found myself unwittingly squeezing my co-laborers’ hands in it, mistaking their hands for the gentle globules. Such an abounding, affectionate, friendly, loving feeling did this avocation beget; that at last I was continually squeezing their hands, and looking up into their eyes sentimentally; as much as to say,—Oh! my dear fellow beings, why should we longer cherish any social acerbities, or know the slightest ill-humor or envy! Come; let us squeeze hands all round; nay, let us all squeeze ourselves into each other; let us squeeze ourselves universally into the very milk and sperm of kindness.

Would that I could keep squeezing that sperm for ever! For now, since by many prolonged, repeated experiences, I have perceived that in all cases man must eventually lower, or at least shift, his conceit of attainable felicity; not placing it anywhere in the intellect or the fancy; but in the wife, the heart, the bed, the table, the saddle, the fire-side, the country; now that I have perceived all this, I am ready to squeeze case eternally. In thoughts of the visions of the night, I saw long rows of angels in paradise, each with his hands in a jar of spermaceti.

And this paragraph should immediately follow the Melville. Now, we check \includedoc and \includedoc* in a similar manner. The included file is going to use the alltt environment. We have loaded the package in the parent file. The \usepackage command in the included file should be disabled; if not, there will be an error when it tries to redefine the environment. Also, we disable several commands that will produce an error if they are not. This should be the last text on this page.

This line is due to the preamble of the included file. This file could be compiled on its own.

Here is yet another quotation from Beckett, this time from Watt
(45--46).

The crocuses and the larch turning green every year a week before the others and the pastures red with uneaten sheep’s placentas and the long summer days and the new-mown hay and the wood-pigeon in the morning and the cuckoo in the afternoon and the corncrake in the evening and the wasps in the jam and the smell of the gorse and the look of the gorse and the apples falling and the children walking in the dead leaves and the larch turning brown a week before the others and the chestnuts falling and the howling winds and the sea breaking over the pier and the first fires and the hooves on the road and the consumptive postman whistling The Roses Are Blooming in Picardy and the standard oil-lamp and of course the snow and to be sure the sleet and bless your heart the slush and every fourth year the February dèbâcle and the endless April showers and the crocuses and then the whole bloody business starting all over again.

Notice, just as one single thing among dozens, that “crocus” sounds like “croak us.”

And this should be on a new page. The first line should have said it comes from the preamble.

The \includedoc family of commands puts the included file into a group. So \preamblecmd is not defined now.

Now we do the same thing over, but the file should appear immediately. This line is due to the preamble of the included file. This file could be compiled on its own.

Here is yet another quotation from Beckett, this time from Watt
(45--46).

The crocuses and the larch turning green every year a week before the others and the pastures red with uneaten sheep’s placentas and the long summer days and the new-mown hay and the wood-pigeon in the morning and the cuckoo in the afternoon and the corncrake in the evening and the wasps in the jam and the smell of the gorse and the look of the gorse and the apples falling and the children walking in the dead leaves and the larch turning brown a week before the others and the chestnuts falling and the howling winds and the sea breaking over the pier and the first fires and the hooves on the road and the consumptive postman whistling The Roses Are Blooming in Picardy and the standard oil-lamp and of course the snow and to be sure the sleet and bless your heart the slush and every fourth year the February dèbâcle and the endless April showers and the crocuses and then the whole bloody business starting all over again.

Notice, just as one single thing among dozens, that “crocus” sounds like “croak us.”

And this paragraph should immediately follow the Beckett. Finally, we have to check \includedocskip and \includedocskip*. The difference will be that the first line will say it comes from the parent. This should be the last text on the page.

This def of preamblecmd comes from the parent. This file could be compiled on its own.

Here is yet another quotation from Beckett, this time from Watt
(45--46).

The crocuses and the larch turning green every year a week before the others and the pastures red with uneaten sheep’s placentas and the long summer days and the new-mown hay and the wood-pigeon in the morning and the cuckoo in the afternoon and the corncrake in the evening and the wasps in the jam and the smell of the gorse and the look of the gorse and the apples falling and the children walking in the dead leaves and the larch turning brown a week before the others and the chestnuts falling and the howling winds and the sea breaking over the pier and the first fires and the hooves on the road and the consumptive postman whistling The Roses Are Blooming in Picardy and the standard oil-lamp and of course the snow and to be sure the sleet and bless your heart the slush and every fourth year the February dèbâcle and the endless April showers and the crocuses and then the whole bloody business starting all over again.

Notice, just as one single thing among dozens, that “crocus” sounds like “croak us.”

This should be on a new page. And once again, the quote should follow immediately here:

This def of preamblecmd comes from the parent. This file could be compiled on its own.

Here is yet another quotation from Beckett, this time from Watt
(45--46).

The crocuses and the larch turning green every year a week before the others and the pastures red with uneaten sheep’s placentas and the long summer days and the new-mown hay and the wood-pigeon in the morning and the cuckoo in the afternoon and the corncrake in the evening and the wasps in the jam and the smell of the gorse and the look of the gorse and the apples falling and the children walking in the dead leaves and the larch turning brown a week before the others and the chestnuts falling and the howling winds and the sea breaking over the pier and the first fires and the hooves on the road and the consumptive postman whistling The Roses Are Blooming in Picardy and the standard oil-lamp and of course the snow and to be sure the sleet and bless your heart the slush and every fourth year the February dèbâcle and the endless April showers and the crocuses and then the whole bloody business starting all over again.

Notice, just as one single thing among dozens, that “crocus” sounds like “croak us.”

And this text should immediately follow the quote. END OF TEST